# ManipBench: Benchmarking Vision-Language Models for Low-Level Robot Manipulation

### v1

@@ 2 Related Work
Low-level Manip. Reasoning Manip. Task Factorization Perf. Trans. Eval. Real World Data Fabric Manipulation Dynamic Manipulation Model Diversity
@@
Table 1: Comparison of ManipBench and other vision-language benchmarks or datasets (denoted with *). ManipBench is the only vision-language benchmark that evaluates low-level manipulation reasoning in VLMs using multiple-choice questions, enabling efficient and effective assessment. See the Appendix for more details.
@@ 4 ManipBench: Data Preparation
From public robotic manipulation datasets. We use demonstrations from the open-source robot manipulation datasets of DROID [68] and Bridge [69] for curating ManipBench questions. We study these datasets because of their importance in large-scale robotic imitation learning [48]. We separate the DROID data into two subsets based on the manipulation task: DROID articulated (art.) and DROID pick-and-place (p&p). To prepare the data and generate questions, we extract the ground-truth gripper trajectories to obtain the picking and placing points by finetuning GroundingDINO [70] and leveraging the data from [71], which contains human-annotated gripper positions for affordance prediction across multiple datasets. See the Appendix for details.
From in-house fabric manipulation setup. Given its practical importance [35, 72, 73], we manually curate data specifically designed to evaluate VLMs’ understanding of fabric manipulation, such as folding or smoothing tasks. We break down common aspects of fabric manipulation into ten distinct dimensions, such as understanding of fabric-object interactions and understanding of inverse dynamics. Each dimension represents a fundamental aspect that an agent must implicitly grasp to successfully perform fabric manipulation. We get data from our real-world workstation. This comprises of a 3cm thick 98cm $\times$ 78cm foam, an Intel Realsense d415i RGBD camera mounted at a height of 87.2cm, rectangular fabrics, and solid objects of varying dimensions. This setup is used to capture top-down image observations of numerous scenes in different task settings for formulating questions (Sec. 5.3). The data collection procedure varies slightly across different dimensions. See the Appendix for the description, choice, and importance of the dimensions.
From simulation. To assess manipulation tasks where real-world deployment may be cumbersome, such as tool use and dynamic manipulation, we compile a suite of simulated tasks that spans all task categories in ManipBench: (i) pick-and-place, (ii) articulated object manipulation, (iii) deformable object manipulation, (iv) tool manipulation, and (v) dynamic manipulation. We primarily adapt simulation assets, and pre-trained policies from popular existing benchmarks: SimplerEnv [74], RLBench [67], and SoftGym [59] to generate data for our evaluation questions (Sec. 5.4). We use environments from SimplerEnv for tasks (i) and (ii), while using SoftGym and RLBench, respectively, for tasks (iii) and (iv). For task (v), we construct a new ball-shooting environment in IsaacSim [75]. Additional details on data collection for each task are presented in the Appendix.
@@ 5.2 MCQs from Public Robotic Manipulation Datasets
Overall, we have 9180 questions with 6120 type 1 questions and 3060 type 2 questions. Pure random guessing on type 1 questions will lead to a success rate of 25% since there are always 4 choices, while random guessing for type 2 questions will be worse than 25% performance since it combines predictions from picking and placing. We perform “question augmentation” to build several questions from the same episode. See the Appendix for more details.
@@ 5.3 MCQs from In-house Fabric Manipulation Setup
We pre-process all the image observations following Sec. 5.1. We discard data where the preprocessing pipeline failed (e.g., Grounded SAM could not detect keypoints). With the desired keypoints and grids, we manually create choices for the multiple-choice questions. We generate 2662 questions across all the ten dimensions. We do not perform question augmentation, as done for the other question categories to encourage diverse scenarios being considered, given the simplicity of many dimensions. Thus, in general, we only form one question for each recorded scene. The questions vary across the dimensions, based on what they are trying to evaluate. Pure random guessing will result in 25% success given 4 choices. Additional details and sample questions are in the Appendix.
@@ 5.4 MCQs from Simulation
For the questions generated from existing simulation environments, we do not need to perform pre-processing steps (from Sec. 5.1) given the access to the ground-truth keypoint information. In this case, we plot the ground-truth point, along with sampling some incorrect points, on the image (Fig. 1, the third example). We do not overlay a grid pattern for these questions. We consider candidate object contact points and after-contact movements on the frames to generate evaluation questions. VLMs are required to select one of four candidate keypoints to complete the given manipulation tasks. Pure random guessing will result in 25% success for all tasks. See the Appendix for details.
@@ 6 ManipBench: Evaluation Plan
We evaluate 10 VLM families on ManipBench (see the Appendix for details). This includes GPT and Gemini, two popular industry-backed closed-source model families widely regarded as among the most effective and versatile models in the Generative AI community. The other models are open-source and are known for their effective performance in multi-modal tasks (e.g., InternVL [79] and Qwen-VL [80]). These evaluations were conducted mainly on 2 RTX 4090 GPUs. Another server with 5 RTX 6000 Ada GPUs was used for larger models. Our evaluation metric is the percentage accuracy of correctly answering the questions. For each model and question type, we average over all the questions to report the numbers. As an upper bound baseline to compare with human intuition, we have 36 human volunteers evaluate these questions, through a custom web portal. The volunteers answer a subset of 50-70 questions for each question type, and we use their accuracies to obtain the final human score. We provide more details on the human evaluation in the Appendix.
Real-World Robot Manipulation. We also conduct physical robot manipulation experiments to verify whether performance on ManipBench translates to unseen robot tasks in real world. We design 7 manipulation tasks whose specific combinations of task goals, language instructions, camera angles, and objects never appear together in ManipBench: Color Specific p&p, Spatial Specific p&p, Fit Apple, Pick Crumpled Cloth, Move Fabric Easy, Move Fabric Intermediate, and Move Fabric Hard (Fig. 3) [^1] The starting states for Move Fabric Easy and Move Fabric Hard are the same, so we only show one.. In these experiments, VLMs select keypoints to define robot actions. The pipeline for generating candidate keypoints for real world tasks is similar to the one for generating MCQs in ManipBench (Sec. 5.1). We use a UR5 robot arm with a Robotiq 2F-85 parallel-jaw gripper and a top-down RGBD camera. See the Appendix for more details about the real world experiments.
@@ 7 ManipBench: Results, Analysis, and Key Insights
Model Bridge DROID (art.) DROID (p&p)
@@
Model Place Close Straight. Sweep Ball
@@
Results on MCQs from public robotic manipulation datasets. Table 3 reports the performance of VLMs, random guessing, and human evaluation on the questions from public robotic manipulation datasets (Sec. 5.2). We use red to represent the highest accuracy, orange for the second‑highest, and yellow for the third‑highest, and we repeat this color code for subsequent tables. The Gemini-2.5-pro model attains the best performance across all three type 1 question categories, while o1 performs the best at type 2 questions for DROID (p&p). Among open-source VLMs, InternVL2.5-38B achieves the overall best performance across most questions. While the closed-source VLMs demonstrate a higher-than-random performance, all open-source VLMs with a size smaller than 2B perform similarly to random guessing, if not worse.
Results on MCQs from in-house fabric manipulation setup. See Fig. 4 for a comparison of various VLMs’ performance across multiple dimensions. Overall, VLMs outperform random guessing, suggesting a notable presence of physical reasoning capabilities in the context of fabric manipulation. Certain dimensions, like Task Planning Understanding, are easier for all models, while others, like Fabric-Fabric Interaction Understanding, are more challenging. However, consistent human accuracies across most dimensions suggest that the difficulty is not inherent, revealing gaps in VLMs’ physical reasoning abilities. Among models, Gemini-2.5-pro consistently outperforms others, with o1 being a close contender. We also observe that InternVL2.5-78B usually outperforms the smaller open-source models, though it trails behind Gemini-2.5-pro and o1. Furthermore, the high standard deviations of the model accuracies across different dimensions indicate their effectiveness in distinguishing the models’ low-level inference abilities, particularly in dimensions like Temporal Understanding of Action Sequence and Spatial Reasoning Abilities, justifying our choice of dimensions. Performance of all models on these questions is detailed in the Appendix.
@@ 7.1 Performance Transfer to Unseen Real World Robot Experiments
To demonstrate the utility of ManipBench, we evaluate a set of selected VLMs on 7 unseen real world manipulation tasks (Sec. 6). We perform statistical analysis to assess the correlation between VLM performance on our MCQ benchmark and their effectiveness as real-world robotic agents (Table. 4). On aggregating the accuracies over the different question-types, we observe a Pearson’s coefficient of 0.889 ($p=0.003$), a Spearman’s coefficient of 0.850 ($p=0.007$), and a Kendall’s Tau of 0.691 ($p=0.018$). These results indicate a strong positive correlation between MCQ benchmark performance and real-world effectiveness. The statistically significant ($p<0.05$) Pearson’s coefficient highlights strong linear alignment, while the Spearman and Kendall coefficients confirm robust monotonic agreement with high confidence in rank-order consistency. Collectively, these trends support the utility of ManipBench as a reliable proxy for evaluating VLMs in embodied robotic settings. See the Appendix for details on correlations for each question type with the real-world experiments.
@@ 9 Limitations
While we believe ManipBench is a valuable benchmark for VLMs and robotics, there are some limitations that suggest opportunities for future work. First, due to limited computational and financial resources, we did not conduct experiments on all possible closed-source and open-source model variants, which may draw an incomplete conclusion. Second, MCQ evaluations depend on access to potential correct/incorrect options to choose from. In real-world scenarios where VLMs are employed as agents, they might not always be able to choose among pre-selected options. Third, the benchmark is not exhaustive across different tasks and excludes crucial deformable manipulation tasks that can require low-level reasoning, such as robotic bag manipulation [81, 82, 83]. Finally, ManipBench does not put much emphasis on higher-level aspects (e.g., task planning) since it is focused on low-level reasoning.
@@ Appendix F Additional Results from Public Robotic Manipulation Datasets
Model Bridge DROID (art.) DROID (p&p)
@@
Error modes analysis.
@@ Appendix H Additional Results from the Statistical Analysis
Task Type Pearson Spearman Kendall’s Tau

### v2

@@ 2 Related Work
Low-level
Manip. Reasoning
Manip. Task
Factorization
Perf. Trans.
Eval.
Real World
Data
Fabric
Manipulation
Dynamic
Manipulation
Model
Diversity
@@
Table 1: Comparison of ManipBench and other vision-language benchmarks or datasets (denoted with *). ManipBench is the only vision-language benchmark that evaluates low-level manipulation reasoning in VLMs using multiple-choice questions, enabling efficient and effective assessment. See Appendix A for more details.
@@ 4 ManipBench: Data Preparation
From public robotic manipulation datasets. We use demonstrations from the open-source robot manipulation datasets of DROID [68] and Bridge [69] for curating ManipBench questions. We study these datasets because of their importance in large-scale robotic imitation learning [48]. We separate the DROID data into two subsets based on the manipulation task: DROID articulated (art.) and DROID pick-and-place (p&p). To prepare the data and generate questions, we extract the ground-truth gripper trajectories to obtain the picking and placing points by finetuning GroundingDINO [70] and leveraging the data from [71], which contains human-annotated gripper positions for affordance prediction across multiple datasets. See Appendix B.1 for details.
From in-house fabric manipulation setup. Given its practical importance [35, 72, 73], we manually curate data specifically designed to evaluate VLMs’ understanding of fabric manipulation, such as folding or smoothing tasks. We break down common aspects of fabric manipulation into ten distinct dimensions, such as understanding of fabric-object interactions and understanding of inverse dynamics. Each dimension represents a fundamental aspect that an agent must implicitly grasp to successfully perform fabric manipulation. We get data from our real-world workstation. This comprises of a 3cm thick 98cm $\times$ 78cm foam, an Intel Realsense d415i RGBD camera mounted at a height of 87.2cm, rectangular fabrics, and solid objects of varying dimensions. This setup is used to capture top-down image observations of numerous scenes in different task settings for formulating questions (Sec. 5.3). The data collection procedure varies slightly across different dimensions. See Appendix B.2 for the description, choice, and importance of the dimensions.
From simulation. To assess manipulation tasks where real-world deployment may be cumbersome, such as tool use and dynamic manipulation, we compile a suite of simulated tasks that spans all task categories in ManipBench: (i) pick-and-place, (ii) articulated object manipulation, (iii) deformable object manipulation, (iv) tool manipulation, and (v) dynamic manipulation. We primarily adapt simulation assets, and pre-trained policies from popular existing benchmarks: SimplerEnv [74], RLBench [67], and SoftGym [59] to generate data for our evaluation questions (Sec. 5.4). We use environments from SimplerEnv for tasks (i) and (ii), while using SoftGym and RLBench, respectively, for tasks (iii) and (iv). For task (v), we construct a new ball-shooting environment in IsaacSim [75]. Additional details on data collection for each task are presented in Appendix B.3.
@@ 5.2 MCQs from Public Robotic Manipulation Datasets
Overall, we have 9180 questions with 6120 type 1 questions and 3060 type 2 questions. Pure random guessing on type 1 questions will lead to a success rate of 25% since there are always 4 choices, while random guessing for type 2 questions will be worse than 25% performance since it combines predictions from picking and placing. We perform “question augmentation” to build several questions from the same episode. See Appendix C.1 for more details.
@@ 5.3 MCQs from In-house Fabric Manipulation Setup
We pre-process all the image observations following Sec. 5.1. We discard data where the preprocessing pipeline failed (e.g., Grounded SAM could not detect keypoints). With the desired keypoints and grids, we manually create choices for the multiple-choice questions. We generate 2662 questions across all the ten dimensions. We do not perform question augmentation, as done for the other question categories to encourage diverse scenarios being considered, given the simplicity of many dimensions. Thus, in general, we only form one question for each recorded scene. The questions vary across the dimensions, based on what they are trying to evaluate. Pure random guessing will result in 25% success given 4 choices. Additional details and sample questions are in Appendix C.2.
@@ 5.4 MCQs from Simulation
For the questions generated from existing simulation environments, we do not need to perform pre-processing steps (from Sec. 5.1) given the access to the ground-truth keypoint information. In this case, we plot the ground-truth point, along with sampling some incorrect points, on the image (Fig. 1, the third example). We do not overlay a grid pattern for these questions. We consider candidate object contact points and after-contact movements on the frames to generate evaluation questions. VLMs are required to select one of four candidate keypoints to complete the given manipulation tasks. Pure random guessing will result in 25% success for all tasks. See Appendix C.3 for details.
@@ 6 ManipBench: Evaluation Plan
We evaluate 10 VLM families on ManipBench (see Table 7 for details). This includes GPT and Gemini, two popular industry-backed closed-source model families widely regarded as among the most effective and versatile models in the Generative AI community. The other models are open-source and are known for their effective performance in multi-modal tasks (e.g., InternVL [79] and Qwen-VL [80]). These evaluations were conducted mainly on 2 RTX 4090 GPUs. Another server with 5 RTX 6000 Ada GPUs was used for larger models. Our evaluation metric is the percentage accuracy of correctly answering the questions. For each model and question type, we average over all the questions to report the numbers. As an upper bound baseline to compare with human intuition, we have 36 human volunteers evaluate these questions, through a custom web portal. The volunteers answer a subset of 50-70 questions for each question type, and we use their accuracies to obtain the final human score. We provide more details on the human evaluation in Appendix D.
Real-World Robot Manipulation. We also conduct physical robot manipulation experiments to verify whether performance on ManipBench translates to unseen robot tasks in real world. We design 7 manipulation tasks whose specific combinations of task goals, language instructions, camera angles, and objects never appear together in ManipBench: Color Specific p&p, Spatial Specific p&p, Fit Apple, Pick Crumpled Cloth, Move Fabric Easy, Move Fabric Intermediate, and Move Fabric Hard (Fig. 3) [^1] The starting states for Move Fabric Easy and Move Fabric Hard are the same, so we only show one.. In these experiments, VLMs select keypoints to define robot actions. The pipeline for generating candidate keypoints for real world tasks is similar to the one for generating MCQs in ManipBench (Sec. 5.1). We use a UR5 robot arm with a Robotiq 2F-85 parallel-jaw gripper and a top-down RGBD camera. See Appendix E for more details about the real world experiments.
@@ 7 ManipBench: Results, Analysis, and Key Insights
Model
Bridge DROID (art.) DROID (p&p)
@@
Model
Place Close Straight. Sweep Ball
@@
Results on MCQs from public robotic manipulation datasets. Table 3 reports the performance of VLMs, random guessing, and human evaluation on the questions from public robotic manipulation datasets (Sec. 5.2). We use red to represent the highest accuracy, orange for the second‑highest, and yellow for the third‑highest, and we repeat this color code for subsequent tables. The Gemini-2.5-pro model attains the best performance across all three type 1 question categories, while o1 performs the best at type 2 questions for DROID (p&p). Among open-source VLMs, InternVL2.5-38B achieves the overall best performance across most questions. While the closed-source VLMs demonstrate a higher-than-random performance, all open-source VLMs with a size smaller than 2B perform similarly to random guessing, if not worse. See Appendix F for additional details.
Results on MCQs from in-house fabric manipulation setup. See Fig. 4 for a comparison of various VLMs’ performance across multiple dimensions. Overall, VLMs outperform random guessing, suggesting a notable presence of physical reasoning capabilities in the context of fabric manipulation. Certain dimensions, like Task Planning Understanding, are easier for all models, while others, like Fabric-Fabric Interaction Understanding, are more challenging. However, consistent human accuracies across most dimensions suggest that the difficulty is not inherent, revealing gaps in VLMs’ physical reasoning abilities. Among models, Gemini-2.5-pro consistently outperforms others, with o1 being a close contender. We also observe that InternVL2.5-78B usually outperforms the smaller open-source models, though it trails behind Gemini-2.5-pro and o1. Furthermore, the high standard deviations of the model accuracies across different dimensions indicate their effectiveness in distinguishing the models’ low-level inference abilities, particularly in dimensions like Temporal Understanding of Action Sequence and Spatial Reasoning Abilities, justifying our choice of dimensions. Performance of all models on these questions is detailed in Appendix G.
@@ 7.1 Performance Transfer to Unseen Real World Robot Experiments
To demonstrate the utility of ManipBench, we evaluate a set of selected VLMs on 7 unseen real world manipulation tasks (Sec. 6). We perform statistical analysis to assess the correlation between VLM performance on our MCQ benchmark and their effectiveness as real-world robotic agents (Table. 4). On aggregating the accuracies over the different question-types, we observe a Pearson’s coefficient of 0.889 ($p=0.003$), a Spearman’s coefficient of 0.850 ($p=0.007$), and a Kendall’s Tau of 0.691 ($p=0.018$). These results indicate a strong positive correlation between MCQ benchmark performance and real-world effectiveness. The statistically significant ($p<0.05$) Pearson’s coefficient highlights strong linear alignment, while the Spearman and Kendall coefficients confirm robust monotonic agreement with high confidence in rank-order consistency. Collectively, these trends support the utility of ManipBench as a reliable proxy for evaluating VLMs in embodied robotic settings. See Appendix H for details on correlations for each question type with the real-world experiments.
@@ 9 Limitations
While we believe ManipBench is a valuable benchmark for VLMs and robotics, there are some limitations that suggest opportunities for future work. First, due to limited computational and financial resources, we did not conduct experiments on all possible closed-source and open-source model variants, which may draw an incomplete conclusion. Second, MCQ evaluations depend on access to potential correct/incorrect options to choose from. In real-world scenarios where VLMs are employed as agents, they might not always be able to choose among pre-selected options. Third, the benchmark is not exhaustive across different tasks and excludes crucial deformable manipulation tasks that can require low-level reasoning, such as robotic bag manipulation [81, 82, 83]. Fourth, although we demonstrate a strong correlation between benchmark performance and real-world experiment performance, our current setup leverages VLMs in the same way across both settings where the VLMs select an answer or action from a predefined list of options. A natural next step would have been to explore alternative approaches for real-world experiments, such as requiring VLMs to generate actions directly rather than choosing from fixed options. Finally, ManipBench lacks emphasis on higher-level aspects (e.g., task planning) since it is focused on low-level reasoning.
@@ Appendix D Additional Details about Human Evaluation
On top of the results described in Sec. 6, we also perform a small scale human evaluation for the questions created from simulation environments. Specifically, we ask human volunteers to answer 10 random questions for the Place Carrot and Sweep Object tasks, and observe an accuracy of 100% for both the cases. Due to the relatively smaller scale of these experiments, we do not include them as part of the discussion in the main paper.
@@ Appendix F Additional Results from Public Robotic Manipulation Datasets
Model
Bridge DROID (art.) DROID (p&p)
@@
Error Mode Analysis.
@@
Question Length Analysis.
Our question/answer pairs are very long and they require the model to associate between the text and the image many times. To investigate, we create a small sample of 100 binary multiple choice questions to simplify the stated text-image association. These questions are designed from the Bridge [69] dataset and have the images annotated with two differently colored points (red and blue), with one of the points being a correct answer for the question under consideration.
Each question has two sub-questions with different color palettes for the points. In sub-question 1, the red point is the correct picking point; In sub-question 2, the blue point is the correct picking point. The position of those 2 candidate picking points are identical on both sub-questions. This is done to mitigate possible color bias exhibited by the VLMs while answering these questions. See Fig. 11 for a sample question, and Table 9 for the results. We see that the results with these questions are consistent with the results stated in Section 7. For example, the best VLM is still Gemini-2.5-pro with the overall accuracy of 0.98, while the smaller open-source VLMs demonstrate color bias and thus have relatively poor performance overall.
We also include models of the PaliGemma family [98, 99, 100] in this smaller experimental study. Specifically, the original PaliGemma model with 3 billion parameters (Paligemma-V1-3B), the PaliGemma V2 model with 3 billion parameters (Paligemma-V2-3B), and the Gemma 3 model with 4 billion parameters (Gemma-3-4B). We observe that Gemma 3 performs similar to open-source models of similar size, whereas PaliGemma V2 refuses to answer the questions, resulting in a 0.00 accuracy. Since these models are not evaluated on any other questions in ManipBench, we do not include them in Table 7.
Figure 11: Sample question for performing question length analysis. The task here is “Put the banana on the green cloth,” and the correct choice would be the red point.
Model Accuracy Accuracy Accuracy
Red Correct Blue Correct Both Correct
Closed-Source
Gemini-2.5-pro 1.00 0.98 0.98
GPT-4o 0.60 0.76 0.50
Open-Source
InternVL2-1B 0.34 0.20 0.08
InternVL2-2B 1.00 0.00 0.00
InternVL2-4B 0.12 1.00 0.12
InternVL2-8B 0.22 0.82 0.12
InternVL2-26B 0.56 0.66 0.26
InternVL2.5-1B 0.92 0.00 0.00
InternVL2.5-2B 1.00 0.12 0.12
InternVL2.5-4B 0.10 0.98 0.10
InternVL2.5-8B 0.48 0.74 0.28
InternVL2.5-26B 0.72 0.76 0.48
Qwen2VL-2B 1.00 0.12 0.12
Qwen2VL-7B 0.98 0.58 0.56
Qwen2.5-VL-3B 0.31 0.82 0.22
Qwen2.5-VL-7B 0.51 0.94 0.49
Paligemma-V1-3B 0.76 0.36 0.26
Paligemma-V2-3B 0.00 0.00 0.00
Gemma-3-4B 0.88 0.37 0.32
Random 0.50 0.50 0.25
Table 9: Performance comparison of various VLMs on the simpler binary MCQs, which comprise of red and blue points annotated on the image with one of them being the correct answer for the task in hand. We report the question answering accuracies for the case where the red point is the correct answer, and where the blue point is the correct answer in the first two columns respectively. We also take an intersection of these and report the accuracies of solving both the sub-questions correctly in the third column.
@@ Appendix H Additional Results from the Statistical Analysis
Task Type
Pearson Spearman Kendall’s Tau